# Dual Attention Residual U-Net for Accurate Brain Ultrasound Segmentation in IVH Detection

### v1

@@ 3.4 Loss Function
where I represents the ground truth label , $\hat{\textbf{{I}}}$ represents the predicted probability of the foreground.
Focal loss is used to address class imbalance in segmentation tasks and provide more focus on difficult cases. It assigns higher weights to misclassified pixels, while reducing the impact of easy pixels during training. The Focal loss is computed as follows
@@
In Eq. 9, $\alpha$ is a balancing parameter to control the contribution of each class, and $\gamma$ is the focusing parameter that modulates the rate at which the loss focuses on hard classification cases.
In training, we use a combination of the Dice loss, BCE loss, and Focal loss as shown in Eq. 10 to guide the optimization process. By minimizing this combined loss during training, our model learns to accurately segment the desired regions of interest.
$l=l_{Dice}+l_{BCE}+l_{Focal}.$ (10)
@@ 4.1 Dataset Description
We use the brain anatomy segmentation (Brain US) [24], [27] dataset to evaluate our model. Intraventricular hemorrhage (IVH) resulting in enlargement of the brain ventricles is one of the main causes of preterm brain injury. The main imaging modality used for the diagnosis of brain disorders in preterm newborns is cranial ultrasound because of its safety and cost-effectiveness. In addition, the absence of septum pellucidum is an important biomarker for the diagnosis of septo-optic dysplasia. Automatic segmentation of the brain ventricles and septum pellucidum from these US scans is essential for an accurate diagnosis and prognosis of these diseases. After obtaining institutional review board (IRB) approval, US scans were taken from 20 different premature neonates (age ¡ 1 year). The total number of images collected was 1629 with annotations out of which 1300 were allocated for training and 329 for testing. We resize the images to 128×128 for all our experiments.
@@ 4.2 Evaluation Metrics
The performance of our segmentation model is evaluated using commonly used metrics: dice score (F1) and intersection over union (IoU) score. These metrics provide quantitative measures of the model’s ability to accurately delineate regions of interest.
Intersection over Union
IoU is a measure that quantifies the overlap between the predicted binary segmentation mask and the ground truth mask. It is calculated as the ratio of the intersection area between the two masks to their union area and is defined as follows
@@
Dice Score
The Dice score, also referred to as the F1 score, integrates both recall and precision into a single value for evaluation and is defined as follows
@@ 4.3 Experimental Setup
We have developed our segmentation model using Python and have used the TensorFlow and Keras libraries for implementation. To speed up training, we used the high performance NVIDIA RTX 4060Ti 16G.
@@ 4.4 Hyperparameter Setup
Our model is trained for 150 epochs for the Brain US dataset, where every epoch represents one complete pass through the entire datasets. For the image size, we resize all images to a uniform size of 128×128 pixels, which are input into the model for segmentation. During the training phase, we use the Adam optimizer with a learning rate of 0.0001 to optimize the model. This choice of optimizer allows us to efficiently update the model’s parameters and enhance convergence during training. To ensure a comprehensive assessment, we have divided the data into a split 70% (train), 10% (test) and 20% (validation). We have leveraged the training set to optimize the model’s parameters, fine-tune hyperparameters using the validation set, and gauge the model’s ability to generalize to new data by utilizing the testing set.
4.5 Ablation Study
Upgrading Model from Scratch
@@ 4.5 Ablation Study
    Proposed Res-UNet model with the strengths of CBAM and SAL.
The results in Table 2 show the efficacy of each modification in the upgrading way. Each model has been trained using the linear combination of Dice, BCE, and Focal loss. The CBAM improves performance, while the SAL further enhances accuracy and segmentation quality. Both CBAM and SAL increase Dice and IoU by around 3 points with respect to the baseline. Furthermore, compared to the pure Res-UNet case, the dual attention modules between encoders and decoders contribute about 1.5 to 2 point improvement in terms of both metrics.
Figure 3: Results of the ablation study indicate the improvement in model performance with each experimental modification. GT and PM are Ground Truth and Predicted Mask, respectively. Heatmaps demonstrate the improvement of the model’s performance in focusing on the region of interest.The SAL module filters out the negative impact of query-key matching scores so that the heatmap looks monotonic in the uninformative regions. Note that there is no attention map for case i and case ii.
Figure 3 visually illustrates the enhancement achieved with attention mechanisms. The combination of CBAM and SAL results in improved performance for both large and small regions of interest, refining feature representations, and capturing both global and local spatial dependencies for accurate segmentation. Compared with the CBAM-only case (iii), the full model (iv) performs the segmentation more precisely because the SAL module filters out the ’disturbing’ information from the attention map. Furthermore, the heatmap visualization of the proposed model illustrates the spatial regions where it places more emphasis, showing a close resemblance to the ground truth regions for the Brain US dataset.
Through these experiments and analysis, we are able to improve our segmentation model iteratively, identifying the most effective modifications and attention mechanisms. These advances make a significant contribution to enhancing the accuracy and robustness of our model, positioning it as an advanced solution for segmenting ultrasound images.
Combinations of Loss Functions
We have performed a series of experiments utilizing different loss combinations and have consolidated the results in Table 1. This table highlights the ablation study we have conducted on the loss functions employed to train our model. According to the results in Table 1, we have determined the most efficient loss function during model training and selected the optimal model configuration. Our findings reveal that the combination of Dice, BCE, and focal loss yields the highest performance.
@@ 4.6 State-of-the-art Comparisons
We have performed a comparison between our proposed method and several baseline models along with standard segmentation models. The comparative results, which comprehensively evaluate two metrics, are presented in Table 3. The models compared in the table are well known in the image segmentation field, such as FCN [7], U-Net [16], UNet++ [17], Res-UNet [25], Axial Attention U-Net [19] and Medical Transformer [4]. Our proposed method demonstrates superior performance across these models (see Table 3) in terms of Dice score and IoU, indicating better overall segmentation accuracy. To provide specific performance details, our proposed model achieves a Dice score of 89.04, indicating a higher level of similarity between the ground truth and the predicted segmentation masks. In addition, the IoU metric, with a value of 81.84, indicates the strong capability of the model to accurately delineate regions of interest. Significantly, the proposed method achieves the highest Dice score and IoU out of all the models, suggesting that it excels in terms of segmentation accuracy and overlap with ground truth.
@@
Metrics (i)U-Net[17] (ii)Res-UNet[27] (iii)Res-UNet (w/CBAM only) (iv)Res-UNet (w/CBAM and SAL(Ours))
@@
Overall, the evaluation results show the superior performance of our proposed model compared to the state-of-the-art models. This confirms the effectiveness and robustness of our model in achieving accurate and precise segmentation results, positioning it as a promising solution for various segmentation tasks. Figure 4 shows the segmentation results of our proposed model, demonstrating its ability to segment the regions of interest in ultrasound images accurately. Although the visual result differences between our model and MedT are somehow tiny in most of the masks, it still achieves the best segmentation results by far. This indicates that the model focuses its attention on relevant areas, contributing to its accurate segmentation performance.
Through these comprehensive evaluations and visualizations, our proposed model showcases its potential to significantly improve brain anatomy detection and diagnosis, bringing us closer to the goal of early detection and enhanced patient care.
@@ 4.7 Weakness Analysis
Our model has demonstrated excellent performance in various image segmentation tasks, as depicted in Table 3 and Figure 4. However, it is essential to highlight the failure cases in which non-anatomical regions are misclassified as anatomical ones and vice versa, indicating that the precision and recall are relatively lower. It is important to acknowledge the complexity of the dataset used for the evaluation, which presents challenges in achieving perfect segmentation results. Figure 5 illustrates the difficult cases which our model encounters, resulting in deviations from the ground truth segmentation. These challenges may arise from dataset complexity, variations in image quality, or the presence of ambiguous features that are hard to accurately delineate. Despite these challenges, our proposed model demonstrates a valuable contribution to the field of brain anatomy segmentation.
@@ 4.8 Experimentation on the Nuclei Dataset
To evaluate the effectiveness of our proposed method, we have conducted assessments on another dataset named Nuclei, a well-known collection containing microscopic images of nucleus-stained slides and the corresponding ground truth annotations by expert pathologists. The average resolution for Nuclei images and the corresponding ground truth masks is 2000×2000 pixels. For the evaluation on this dataset, we have maintained consistency using the same set of hyperparameters that are employed in the evaluation of the Brain US dataset. However, the images should be converted to grayscale and resized to the same size as those in the Brain US dataset before being fed to the learning network. The segmentation results of the proposed model on the Nuclei dataset are shown in Figure 6. The visual results show that the overall performance is worse than those for the Brain US dataset because the segmentation regions are numbers of tiny particle-like areas, which make it hard for the attention to focus.
@@ 5 Conclusion and Future Research
The current study presents an innovative segmentation technique called the attention-aided Res-UNet model to detect brain anatomies in ultrasound images. Our model has delivered satisfactory outcomes, revealing its potential to improve detection and diagnosis. Nonetheless, it is crucial to recognize the weakness of our proposed method.
In the weakness analysis section, several potential areas for future research were identified. One particularly promising direction involves investigating more hierarchical architectures that integrate multiple forms of stronger attention modules within the model. Such an approach has the potential to improve performance and deepen our understanding of complex brain anatomy detection challenges using complementary insights. Another direction for future research is to enhance datasets through techniques such as data enhancements, which can increase their diversity and size, thereby enhancing generalization and robustness.
In addition, evaluating the robustness of the model could provide valuable insight. Performing segmentation on other types of medical images, as we did in Nuclei dataset, would be beneficial in assessing the model’s versatility and applicability in diverse medical imaging tasks.

### v2

@@ 3.4 Loss Function Combinations
where I represents the ground truth label , $\hat{\textbf{I}}$ represents the foreground prediction.
Focal loss is used to concentrate on difficult cases in image segmentation tasks. The larger weights are given to the hard part of the image pixels, while the smaller ones are assigned to the easy pixels during training. The Focal loss is formulated as follows
@@
In (8), $\alpha$ is a balancing parameter and $\gamma$ is the focusing parameter.
In training, we tried using a single loss function and also combinations of these functions, as shown in forms of (9) to monitor the optimization process.
$l=\alpha l_{Dice}+\beta l_{BCE}+\lambda l_{Focal}.$ (9)
@@ 4.1 Dataset Description
We use the brain anatomy segmentation data set (Brain US) [27], [30] to evaluate our model. Intraventricular hemorrhage (IVH) resulting in enlargement of the brain ventricles is one of the main causes of preterm brain injury. The main imaging modality used for the diagnosis of brain disorders in preterm newborns is cranial ultrasound because of its safety and cost-effectiveness. In addition, the absence of septum pellucidum is an important biomarker for the diagnosis of septo-optic dysplasia. Automatic segmentation of the brain ventricles and septum pellucidum from these US scans is essential for an accurate diagnosis and prognosis of these diseases. After obtaining institutional review board (IRB) approval, US scans were taken from 20 different premature neonates (age less than 1 year). The total number of images collected was 1629 with annotations out of which 1300 were allocated for training and 329 for testing. We resize the images to 128×128 for all our experiments.
@@ 4.2 Evaluation Metrics
Tow metrics are used to evaluate the performance of our segmentation model: dice score (F1) and intersection over union score (IoU). These metrics provide a quantitative way to evaluate the performance of the segmentation process.
4.2.1 Intersection over Union
IoU is a measure that computes the overlap between the predicted segmentation mask and the ground truth mask. It is formulated as follows
@@ 4.2.1 Intersection over Union
4.2.2 Dice Score
The Dice score, also referred to as the F1 score, is formulated as follows
@@ 4.3 Experimental Setup
We have developed our segmentation model using Python and have used the PyTorch library for implementation. To speed up training, we used the high performance NVIDIA RTX 4060Ti 16G.
@@ 4.4 Hyperparameter Setup
Our model is trained for 150 epochs for the Brain US dataset. For the image size, we resize all images to a uniform size of 128×128 pixels, which are fed into the segmentation model. During the training phase, we use the Adam optimizer with a learning rate of 0.00001 to optimize the model. To ensure a comprehensive evaluation, we use the common split strategy and divide the data into a split: 70% (train), 10% (test) and 20% (validation). All the setups discussed above are usually common in such a training task.
Figure 3: Results of the ablation study indicate the improvement in model performance with each experimental modification. GT and PM are ground truth and predicted mask, respectively. Heatmaps demonstrate the improvement of the model’s performance in focusing on the region of interest.The SAL branch filters out the negative impact of query-key matching scores so that the heatmap looks monotonic in the uninformative regions. Note that there is no attention map for case i and case ii.
4.5 Ablation Tests
4.5.1 Upgrading Model from Scratch
@@
    Proposed Res-UNet model with the strengths of CBAM and HAL.
The results in Table 2 show the efficacy of each modification in the upgrading way. Each model has been trained using the linear combination of Dice, BCE, and Focal loss with average weights. Both CBAM and HAL increase Dice and IoU by around 3 points with respect to the baseline. Furthermore, compared to the pure Res-UNet case, the dual attention modules between encoders and decoders contribute about 1.5 to 2 point improvement in terms of both metrics.
Fig. 3 visually illustrates the performance achieved with our combinations of CBAM and HAL attention mechanisms. The entire framework can refine feature representations and capture both global and local spatial correlations for accurate segmentation. Compared with the CBAM-only case (iii), the full model (iv) performs the segmentation more precisely because the SAL branch filters out the ’disturbing’ information from the attention map. Furthermore, the heatmap visualization of the proposed model illustrates the spatial regions where it places more emphasis, showing a close resemblance to the ground truth regions for the Brain US dataset.
4.5.2 Combinations of Loss Functions
We have performed some incremental tests on different combinations of loss functions and have consolidated the results in Table 1. Based on the results in Table 1, we have compared the Dice and IoU scores for each combination. Our conclusion is that the linear combination of all three loss functions produces the best result in our segmentation tests. Therefore, we applied this paradigm for all the experiments in the paper.
@@ 4.6 State-of-the-art Comparisons
We have performed a comparison between our proposed method and several baseline models along with standard segmentation models. The comparative results, which comprehensively evaluate two metrics, are presented in Table 3. The models compared in the table are well known in the field of image segmentation, such as FCN [7], U-Net [15], UNet++ [16], Res-UNet [28], Axial Attention U-Net [18], and Medical Transformer [4]. Our proposed method demonstrates superior performance across these models (see Table 3) in terms of Dice score and IoU, indicating better overall segmentation accuracy. To provide specific performance details, our proposed model achieves a Dice score of 89.04, indicating a higher level of similarity between the ground truth and the predicted segmentation masks. In addition, the IoU metric, with a value of 81.84, indicates the strong capability of the model to accurately delineate regions of interest. In conclusion, the proposed model achieves the highest Dice scores out of all the models, suggesting that it performs better than other models.
@@
Metrics (i)U-Net[17] (ii)Res-UNet[27] (iii)Res-UNet (w/CBAM only) (iv)Res-UNet (w/CBAM and HAL(Ours))
@@
Fig. 4 shows the segmentation results of our proposed model, demonstrating its ability to segment the regions of interest in ultrasound images accurately. Although the visual result differences between our model and MedT are somehow tiny in most of the masks, it still achieves the best segmentation results by far. This suggests that the model pays attention to only relevant areas, which is beneficial to its segmentation accuracy.
By these comprehensive evaluations and visualizations, our proposed model demonstrates its ability to significantly improve the detection and diagnosis of brain anatomy , leading to the goal of early alarming.
@@ 4.7 Weakness Analysis
Our model has shown superior performance in various image segmentation tasks, as depicted in Table 3 and Fig. 4. However, it is essential to highlight the failure cases in which non-anatomical regions are misclassified as anatomical ones and vice versa, indicating that the precision and recall are relatively lower. Intuitively, the quality of segmentation results is dependent on the complexity of the dataset for training and evaluation. Fig. 5 illustrates the difficulties encountered by our model, leading to deviations from the ground truth. These challenges may arise from the complexity of the dataset, variations in image quality, or the presence of ambiguous features that are difficult to delineate accurately. Despite these challenges, our proposed model has shown valuable contributions in the field of anatomical segmentation.
@@ 4.8 Experimentation on the Nuclei Dataset
In order to evaluate the effectiveness of our proposed method, we performed an evaluation on another dataset named Nuclei, which is a well-known collection that contains bright microscopy images of stained cells and correctly annotated masks by expert pathologists. The average resolution of the Nuclei images and their corresponding correct annotation masks is 2000×2000 pixels. For the evaluation of this dataset, we used the same set of hyperparameters as used in the evaluation of the Brain US data set to maintain consistency. The images should be converted to grayscale and resized to the same size as the images in the Brain US dataset before being fed into the network. The segmentation results of the proposed model on the Nuclei dataset are shown in Fig. 6. The visual results show that the overall performance is worse than the performance of the Brain US dataset, as the segmentation area is made up of many tiny grainy regions, which makes it difficult to focus attention.
@@ 5 Conclusion and Future Research
The current research proposes an innovative segmentation technique, called attention-based Res-UNet model, to detect brain anatomy in ultrasound images. Our model has provided satisfactory results, revealing its potential to improve detection and diagnosis. In the weaknesses section, some potential areas for future research were identified. One particularly promising way involves studying more advanced architectures that integrate multiple more powerful attention within the model. This approach has the potential to improve performance and expand our understanding of the complex challenges in detecting brain anatomy. Another direction for future research involves the dataset using techniques such as data augmentation, which can increase their diversity and size, leading to improved generalization and robustness. Furthermore, generalizing segmentation on other types of medical images, as we did for the nuclei data set, will assess the versatility and applicability of the model to a variety of medical imaging tasks.